The Walton Toy Company manufactures a line of dolls and a sewing kit. Demand for the company’s products is increasing and management requests assistance from you in determining an economical sales and production mix for the coming year. The company has provided the following data:

Product Demand
Next year
(units)
Selling
Price
per Unit
Direct
Materials
Direct
Labor
Debbie 50,000 $ 16.70 $ 4.30 $ 6.40
Trish 42,000 $ 7.50 $ 1.10 $ 4.00
Sarah 35,000 $ 26.60 $ 6.44 $ 11.20
Mike 40,000 $ 14.00 $ 2.00 $ 8.00
Sewing kit 325,000 $ 9.60 $ 3.20 $ 3.20

The following additional information is available:  

  1. The company’s plant has a capacity of 130,000 direct labor-hours per year on a single-shift basis. The company’s present employees and equipment can produce all five products.

  2. The direct labor rate of $16 per hour is expected to remain unchanged during the coming year.

  3. Fixed manufacturing costs total $520,000 per year. Variable overhead costs are $2 per direct labor-hour.

  4. All of the company’s nonmanufacturing costs are fixed.

  5. The company’s finished goods inventory is negligible and can be ignored.

Required:

1. How many direct labor hours are used to manufacture one unit of each of the company’s five products?

2. How much variable overhead cost is incurred to manufacture one unit of each of the company’s five products?

3. What is the contribution margin per direct labor-hour for each of the company’s five products?

4. Assuming that direct labor-hours is the company’s constraining resource, what is the highest total contribution margin that the company can earn if it makes optimal use of its constrained resource?

5. Assuming that the company has made optimal use of its 130,000 direct labor-hours, what is the highest direct labor rate per hour that Walton Toy Company would be willing to pay for additional capacity (that is, for added direct labor time)?
Answer #1
Debbie Trish Sarah Mike Sewing Kit
Demand 50000 42000 35000 40000 325000
Selling price per unit $           16.70 $               7.50 $           26.60 $        14.00 $          9.60
Direct material cost per unit $             4.30 $               1.10 $              6.44 $           2.00 $          3.20
Direct labor cost per unit $             6.40 $               4.00 $           11.20 $           8.00 $          3.20
2) Variable overhead per unit $             0.80 $               0.50 $              1.40 $           1.00 $          0.40
Total variable cost per unit $           11.50 $               5.60 $           19.04 $        11.00 $          6.80
Contribution margin per unit $             5.20 $               1.90 $              7.56 $           3.00 $          2.80
1) Direct labor hours per unit 0.40 0.25 0.70 0.50 0.20
3) CM per direct labor hour $           13.00 $               7.60 $           10.80 $           6.00 $        14.00
Ranking of products 2 4 3 5 1
4) Total direct labor hours required 20000 10500 24500 20000 65000
Allocation of DLH: DLH Total DLH Total CM
Sewing kit 65000 65000 $     9,10,000
Debbie 20000 85000 $     2,60,000
Sarah 24500 109500 $     2,64,600
Trish 10500 120000 $         79,800
Mike 10000 130000 $         60,000
Total contribution margin $   15,74,400
5) Total labor cost that can be paid = CM lost on Mike = 10000*6= $         60,000
Labor hours required = 10000
Highest direct labor hours that the firm can pay = 60000/10000 = $              6.00